This will be an expository talk. We shall first describe some of the guiding problems of arithmetic geometry which concern the structure of rational points on algebraic varieties, and try to give a general picture of their current status. We shall then focus on the existence problem for rational varieties and its relation with Brauer groups. As time permits, we shall mention recent progress in the case of (not necessarily rational) elliptic surfaces, with applications to rational varieties.